# First-return statistics in Henyey–Greenstein scattering: Motzkin polynomials and the Cauchy kernel

### v4

@@ Abstract
We study first-return statistics for photons undergoing three-dimensional Henyey–Greenstein scattering in a semi-infinite medium. In previous work, we showed that one-dimensional first-passage probabilities expand in Catalan and Motzkin generating functions. Extending to three dimensions requires a Boundary Truncation Factor (BTF) that accounts for the restricted angular phase space imposed by the boundary. Extensive Monte Carlo simulations are used to determine the BTF empirically as a function of scattering order and anisotropy. For moderate anisotropy, the BTF is accurately described by a Cauchy kernel with parameters depending only on the Henyey–Greenstein asymmetry factor. This closed-form expression reproduces Monte Carlo results to 1–2% accuracy over a broad range of scattering orders. At higher anisotropy, systematic deviations from the Cauchy form are observed and can be reduced using a one-parameter generalized kernel. The resulting framework provides a computationally efficient mapping from three-dimensional anisotropic transport to one-dimensional combinatorial first-passage theory.
Keywords: first-passage problem, random walk, Cauchy kernel, Motzkin polynomials, radiative transfer, Henyey–Greenstein scattering
@@ 1 Introduction
First-passage problems appear throughout statistical mechanics—in polymer physics, diffusion-limited aggregation, financial mathematics, and queueing theory [1, 2, 3]. A clean example arises in radiative transport: photons entering a scattering medium execute a three-dimensional random walk and may return to the entry boundary. We consider a semi-infinite medium occupying the half-space $z>0$, with the entry boundary at $z=0$; photons enter at normal incidence. The statistics of this first-passage event encode how stochastic motion interacts with geometric constraint.
@@
In our previous work in this journal [6], we established that first-return probabilities in one-dimensional isotropic scattering expand in Catalan numbers, the combinatorial objects counting Dyck paths. The reflectance of a semi-infinite Kubelka–Munk medium [7] admits the generating function representation
$R_{\infty}(S,\chi)=\frac{1}{2}\cdot\frac{S}{S+\chi}\cdot C\left(\frac{S^{2}}{4(S+\chi)^{2}}\right)$ (1)
where $C(x)=\sum_{n=0}^{\infty}C_{n}x^{n}$ is the Catalan generating function [17] and $C_{n}=(2n)!/(n!(n+1)!)$. This is distribution-free: it depends only on the zigzag structure, not step lengths. Forward-peaked scattering introduces “flat” steps—events with no $z$-direction change. The Motzkin extension [14, 15] handles this, giving
$P_{\mathrm{refl}}^{(1D)}(n,r)=M_{n-2}\left(\frac{1-2r}{r}\right)\cdot\left(\frac{r}{2}\right)^{n-1}$ (2)
where $M_{n}(t)$ is the Motzkin polynomial and $r$ the backward-step probability.
@@
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-n_{0}}{m_{x}(g)}\right)^{2}}$ (3)
with parameters expressible in terms of the anisotropy factor $g$ alone: {align} m_x(g) = 4g1-g   $\text$(width)
A(g) = 1 - g(1+g)2   $\text$(amplitude)
n_0 = 2   $\text$(peak location) The simple integer coefficients suggest underlying geometric structure; however, a first-principles derivation of the Cauchy form remains an open problem (Section 7.1).
@@
Paper / print media 0.40.40.60.6 [22]
@@
Human dermis/epidermis 0.70.70.90.9 [20]
Biological tissue (in vivo) 0.90.90.980.98 [21]
Intralipid phantoms $\sim 0.9$ Standard value
@@ 1.1 Relation to previous work
Incidence Normal only Normal only
@@
From our 2020 paper, we carry forward: (i) the Motzkin polynomial framework (equation (2)); (ii) the distribution-free character of first-passage combinatorics; and (iii) the connection to classical fluctuation theory [5, 4]. What is new here is: (i) identification of the Cauchy kernel form for the BTF through systematic model selection; (ii) determination of the parameters $m_{x}(g)$ and $A(g)$ from Monte Carlo fitting; and (iii) a modified Cauchy kernel extending validity to high anisotropy.
@@ 1.2 Comparison with generalized Kubelka–Munk
Our approach differs from previous 3D extensions of Kubelka–Munk. Sandoval and Kim [9, 10] extended KM through double spherical harmonics ($\mathrm{DP}_{1}$), obtaining an $8\times 8$ system for forward and backward power flow. For isotropic scattering in optically thick media, their generalized KM achieves errors below 15% when $z_{0}\geq 10$ (where $z_{0}$ is the optical thickness).
However, $\mathrm{DP}_{1}$ encounters difficulties for anisotropic scattering. Its basis functions contain only first-order azimuthal harmonics and cannot capture forward-peaked phase functions. At $g=0.8$, Sandoval and Kim found errors exceeding 80% in transmitted power; at higher anisotropy, the approximation gives negative intensities.
@@
Angular basis $\mathrm{DP}_{1}$ (4 func./hemisphere) HG sampling + Motzkin
@@
Incidence Oblique via eigenmode Normal (oblique: future work)
@@
Practitioner guidance. For semi-infinite media with $g\lesssim 2/3$, BTF+Motzkin is more accurate than $\mathrm{DP}_{1}$: 1D polynomial evaluation achieves sub-2% accuracy versus 15%+ for gKM. For finite slabs or strong absorption, use gKM or Monte Carlo. BTF excels in inverse problems requiring thousands of transport evaluations.
@@ 1.3 Paper organization
Section 2 reviews the 1D Motzkin framework [6]. Section 3 introduces the Boundary Truncation Factor and its Cauchy kernel form. Section 4 describes the Monte Carlo procedure. Section 5 presents the computational algorithm. Section 6 presents the modified Cauchy kernel for high-anisotropy scattering ($g>2/3$). Section 7 concludes.
@@ 2.1 Kubelka–Munk reflectance and Catalan numbers
[…]
where $I$ and $J$ are forward and backward fluxes, $S$ the scattering coefficient, and $\chi$ the absorption coefficient. The reflectance of a semi-infinite slab is
$R_{\infty}(S,\chi)=\frac{S+\chi}{S}-\sqrt{\left(\frac{S+\chi}{S}\right)^{2}-1}$ (5)
@@
$R_{\infty}(S,\chi)=\sum_{n_{p}=1}^{\infty}\frac{C_{n_{p}-1}}{2^{2n_{p}-1}}\left(\frac{S}{S+\chi}\right)^{2n_{p}-1}$ (6)
@@ 2.2 Motzkin extension for forward scattering
The one-dimensional first-return probability with forward scattering becomes equation (2), encoding three processes: up-steps (positive $z$-projection), down-steps (negative $z$-projection), and flat steps (zero $z$-projection change).
@@ 3.1 Physical origin
In bulk scattering without boundaries, Pfeiffer and Chapman [13] proved that the Henyey–Greenstein phase function is closed under successive scattering: after $n$ scattering events, the angular distribution of the scattering cosine $\cos\theta=\bm{\Omega}_{0}\cdot\bm{\Omega}_{n}$ is itself a Henyey–Greenstein distribution with parameter $g^{n}$:
@@
$g_{\mathrm{eff}}^{\mathrm{(constrained)}}=g^{n}\cdot\mathrm{BTF}(n,g)$ (10)
@@
[…]
@@ 3.2 Empirical discovery
For each $(g,n)$ pair, we extracted the BTF as the value needed to make the 1D Motzkin formula reproduce the Monte Carlo 3D result:
$P_{\mathrm{3D}}^{\mathrm{(MC)}}(n,g)=P_{\mathrm{1D}}^{\mathrm{(Motzkin)}}(n,r_{b}(g,n))$ (12)
@@ 3.3 The Cauchy kernel
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-2}{m_{x}(g)}\right)^{2}}$ (13)
with parameters: {align} A(g) = 1 - g(1+g)2 = (1-g)(2+g)2   $\text$(amplitude)
m_x(g) = 4g1-g   $\text$(width) The peak location $n_{0}=2$ corresponds to the minimum scattering order for first return. This parameterized form reproduces Monte Carlo-derived BTF values with mean deviation <<2% and cross-validated $R^{2}>0.999$ for $g\leq 2/3$.
@@
Figure 1 shows the BTF directly as a function of scattering order $n$ for several anisotropy values. The Cauchy profile is evident: a peak at $n=2$ with power-law decay at large $n$.
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation 13). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible.
@@ 4.1 Simulation procedure
We follow the standard Monte Carlo procedure for photon transport [18]. First return occurs when $z$ first becomes negative [19]. Photons are initialized at the origin ($z=0^{+}$) with direction cosine $\mu_{0}=1$. At each scattering event, path length is sampled from $p(c)=e^{-c}$ (unit mean free path) and scattering angle from the Henyey–Greenstein phase function via inverse CDF. The scattering order $n$ is recorded at first occurrence of $z<0$. We simulated $10^{8}$ trajectories per $g$ value with 10 independent runs.
@@
$P_{\mathrm{HG}}(\mu;g)=\frac{1-g^{2}}{2(1+g^{2}-2g\mu)^{3/2}},\quad-1\leq\mu\leq 1$ (14)
where $g=\langle\mu\rangle\in[0,1)$ is the anisotropy parameter.
@@ 4.2 Fit quality
Error bounds. For $g<2/3$, root-mean-square deviation is below 2% across all $n$. For $g>2/3$, errors grow approximately as $\epsilon(g)\approx 0.05\cdot(g-2/3)/(1-g)$, reaching $\sim$15% at $g=0.9$. For $g>0.8$, the modified kernel (Section 6) is recommended.
5 Computational algorithm
We present the complete algorithm for mapping 3D Henyey–Greenstein scattering to 1D Motzkin combinatorics. The algorithm has four steps: (1) determine the angular threshold separating forward from backward scattering; (2) compute the BTF-modified exponent; (3) evaluate the effective backscattering probability; (4) apply the Motzkin formula.
Step 1 (Angular threshold): The threshold $\mu_{b}(g)$ is the direction cosine separating “effective backward” from “effective forward” scattering in the 1D projection. Physically, $\mu_{b}$ partitions the angular phase space so that the 1D Motzkin counting matches the 3D return statistics. Solve
[…]
where $p_{r2}(g)=\int_{-1}^{0}\frac{\mu^{\prime}-1}{\mu^{\prime}}P_{\mathrm{HG}}(\mu^{\prime},g)\,\mathop{}\!\mathrm{d}\mu^{\prime}$ is the exact two-step return probability, $F(\mu;g)$ is the HG cumulative distribution function, and $g^{2}$ appears because after two scattering events the angular distribution has parameter $g^{2}$ (equation 9).
Step 2 (Effective exponent): In bulk scattering, the effective anisotropy after $n$ events is $g^{n}$ (standard convolution). Near a boundary, transverse filtering reduces this—captured by the BTF. We model the boundary-constrained effective exponent as
$\gamma(g,n)=2+\mathrm{BTF}(n,g)(n-2)$ (16)
This interpolates between $\gamma=2$ at $n=2$ (where BTF =A(g)=A(g) is maximal and the constraint is weakest) and $\gamma\to 2$ as $n\to\infty$ (where BTF $\to 0$ and boundary truncation dominates). The baseline $\gamma=2$ reflects the minimum scattering order for first return. This ansatz is empirically motivated: it produces the correct limits and matches Monte Carlo across the parameter space.
Step 3 (Effective backscattering): Substitute into the HG CDF:
$r_{b}(g,n)=F(-\mu_{b}(g);g^{\gamma(g,n)})$ (17)

### v5

@@ Abstract
We study first-return statistics for photons undergoing three-dimensional Henyey–Greenstein scattering in a semi-infinite medium. In previous work, we showed that one-dimensional first-passage probabilities expand in Catalan and Motzkin generating functions. Extending to three dimensions requires a Boundary Truncation Factor (BTF) that accounts for the restricted angular phase space imposed by the boundary. Extensive Monte Carlo simulations are used to determine the BTF empirically as a function of scattering order and anisotropy. For moderate anisotropy, the BTF is accurately described by a Cauchy kernel with parameters depending only on the Henyey–Greenstein asymmetry factor. This closed-form expression reproduces Monte Carlo results to 1–2% accuracy over a broad range of scattering orders. At higher anisotropy, systematic deviations from the Cauchy form are observed and can be reduced using a one-parameter generalized kernel. We further extend the framework to oblique incidence by replacing the normal-incidence return probability with a Legendre-series formula; the BTF parameters and Motzkin counting machinery are independent of incidence angle, so only the anchor point of the algorithm changes. The resulting framework provides a computationally efficient mapping from three-dimensional anisotropic transport at arbitrary incidence to one-dimensional combinatorial first-passage theory.
Keywords: first-passage problem, random walk, Cauchy kernel, Cauchy BTF conjecture, Motzkin polynomials, radiative transfer, Henyey–Greenstein scattering, oblique incidence
@@ 1 Introduction
First-passage problems appear throughout statistical mechanics—in polymer physics, diffusion-limited aggregation, financial mathematics, and queueing theory [1, 2, 3]. A clean example arises in radiative transport: photons entering a scattering medium execute a three-dimensional random walk and will eventually return to the entry boundary. We consider a semi-infinite medium occupying the half-space $z>0$, with the entry boundary at $z=0$; photons enter as a collimated pencil beam at incidence angle $\theta_{\mathrm{inc}}$ (direction cosine $\mu_{\mathrm{inc}}=\cos\theta_{\mathrm{inc}}$). Physically, this corresponds to the reflectance of a narrow collimated beam incident on a scattering half-space—the probability that the beam is scattered back across the entry surface, resolved by the number of scattering events. The statistics of this first-passage event encode how stochastic motion interacts with geometric constraint.
This is a canonical problem of the scalar radiative transfer equation (RTE). The quantity we compute—the scattering-order-resolved first-return probability $P(n,g,\mu_{\mathrm{inc}})$—is more fundamental than the reflectance itself: it serves as a generating function for reflectance at arbitrary single-scattering albedo $a\in[0,1]$, since $R(g,a)=\sum_{n=2}^{\infty}P(n,g)\,a^{n}$. The entire framework operates at pure scattering ($a=1$); absorption enters only through this weighted sum, which amounts to a discrete Laplace transform. Because $P(n,g)$ depends only on the phase function geometry, a single evaluation of the framework covers all albedos simultaneously. The result is an analytical alternative to Monte Carlo for this geometry: one polynomial evaluation per scattering order, rather than $10^{8}$ photon trajectories.
The practical significance lies in imaging and inverse problems. Diffuse optical tomography, tissue characterization, and remote sensing all require solving the RTE repeatedly—often thousands of times—while fitting scattering parameters to measured reflectance data. When each forward evaluation requires a full Monte Carlo simulation, the computational cost becomes prohibitive. A closed-form forward model that evaluates in microseconds rather than minutes transforms the feasibility of these inverse problems. The same applies to industrial applications such as paper and print characterization [22], where rapid parameter estimation from reflectance measurements is standard practice.
@@
In our previous work [6], we established that first-return probabilities in one-dimensional isotropic scattering expand in Catalan numbers, the combinatorial objects counting Dyck paths. The reflectance of a semi-infinite Kubelka–Munk medium [7] admits the generating function representation
$R_{\infty}(S,K)=\frac{1}{2}\cdot\frac{S}{S+K}\cdot C\!\left(\frac{S^{2}}{4(S+K)^{2}}\right)$ (1)
where $C(x)=\sum_{n=0}^{\infty}C_{n}x^{n}$ is the Catalan generating function [17] and $C_{n}=(2n)!/(n!(n+1)!)$. This is distribution-free: it depends only on the zigzag structure, not step lengths.
Forward-peaked scattering introduces “flat” steps—events with no $z$-direction change. The Motzkin extension [14, 15] handles this. The one-dimensional first-return probability with forward scattering is given by the marginal formula:
$P_{\mathrm{1D}}^{\mathrm{marg}}(n,r)=\sum_{n_{p}=1}^{\lfloor n/2\rfloor}\frac{1}{2^{2n_{p}-1}}\,\frac{(n-2)!}{(n-2n_{p})!\,n_{p}!\,(n_{p}-1)!}\;r^{2n_{p}-1}(1-r)^{n-2n_{p}}$ (2)
where $n$ is the scattering order ($n\geq 2$), $n_{p}$ is the peak count (number of direction reversals), $r$ is the backward-step probability, and (1r)(1-r) is the non-backward probability. The combinatorial coefficient is the Motzkin triangle number $T(n-2,n_{p}-1)$.
@@
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\dfrac{n-n_{0}}{m_{x}(g)}\right)^{2}}$ (3)
with parameters expressible in terms of the anisotropy factor $g$ alone:
$m_{x}(g)$ $=\frac{4g}{1-g}\qquad\text{(width)}$ (4)
$A(g)$ $=1-\frac{g(1+g)}{2}\qquad\text{(amplitude)}$ (5)
$n_{0}$ $=2\qquad\text{(peak location)}$ (6)
The simple integer coefficients suggest underlying geometric structure; however, a first-principles derivation of the Cauchy form remains an open problem (Section 8). We state this as a formal conjecture in Section 3.3.
@@
A key feature of the framework is that the BTF parameters $A(g)$ and $m_{x}(g)$ are independent of incidence angle. This allows a clean extension to oblique incidence (Section 6): only the single-scattering return probability—the anchor point of the algorithm—changes, while the entire Motzkin counting machinery carries through unchanged.
@@
Paper / print media 0.4–0.6 [22]
@@
Human dermis/epidermis 0.7–0.9 [20]
Biological tissue (in vivo) 0.9–0.98 [21]
Intralipid phantoms $\approx 0.9$ Standard value
@@ 1.1 Relation to previous work
From our 2020 paper, we carry forward: (i) the Motzkin polynomial framework (equation (2)); (ii) the distribution-free character of first-passage combinatorics; and (iii) the connection to classical fluctuation theory [4, 5]. What is new here is: (i) identification of the Cauchy kernel form for the BTF through systematic model selection; (ii) determination of the parameters $m_{x}(g)$ and $A(g)$ from Monte Carlo fitting; (iii) a modified Cauchy kernel extending validity to high anisotropy; and (iv) extension to oblique incidence via a Legendre-series formula.
@@
Incidence Normal only Normal and oblique
@@ 1.2 Comparison with generalized Kubelka–Munk
Our approach differs from previous 3D extensions of Kubelka–Munk. Sandoval and Kim [9, 10] extended KM through double spherical harmonics (DP1), obtaining an $8\times 8$ system for forward and backward power flow. For isotropic scattering in optically thick media, their generalized KM achieves errors below 15% when $z_{0}\geq 10$ (where $z_{0}$ is the optical thickness).
However, DP1 encounters difficulties for anisotropic scattering. Its basis functions contain only first-order azimuthal harmonics and cannot capture forward-peaked phase functions. At $g=0.8$, Sandoval and Kim found errors exceeding 80% in transmitted power; at higher anisotropy, the approximation gives negative intensities.
@@
Practitioner guidance. For semi-infinite media with $g\lesssim 2/3$, the BTF framework is more accurate than DP1: 1D polynomial evaluation achieves sub-2% accuracy versus 15%+ for gKM. For finite slabs or strong absorption, use gKM or Monte Carlo. BTF excels where a fast forward model is needed: iterative parameter fitting in diffuse optical tomography, reflectance-based tissue characterization, and any inverse problem requiring thousands of transport evaluations.
@@
Angular basis DP1 (4 func./hemisphere) HG sampling + Motzkin
@@
Incidence Oblique via eigenmode Normal and oblique
@@ 1.3 Paper organization
Section 2 reviews the 1D Motzkin framework [6]. Section 3 introduces the Boundary Truncation Factor and its Cauchy kernel form. Section 4 describes the Monte Carlo procedure. Section 5 presents the computational algorithm for normal incidence. Section 6 extends the framework to oblique incidence. Section 7 presents the modified Cauchy kernel for high-anisotropy scattering ($g>2/3$). Section 8 concludes.
@@ 2.1 Kubelka–Munk reflectance and Catalan numbers
$\begin{pmatrix}dI\\ dJ\end{pmatrix}=\begin{pmatrix}-(S+K)&S\\ -S&(S+K)\end{pmatrix}\begin{pmatrix}I\\ J\end{pmatrix}dz$ (7)
where $I$ and $J$ are forward and backward fluxes, $S$ the scattering coefficient, and $K$ the absorption coefficient. The reflectance of a semi-infinite slab is
$R_{\infty}(S,K)=\frac{S+K}{S}-\sqrt{\left(\frac{S+K}{S}\right)^{2}-1}$ (8)
@@
$R_{\infty}(S,K)=\sum_{n_{p}=1}^{\infty}\frac{C_{n_{p}-1}}{2^{2n_{p}-1}}\left(\frac{S}{S+K}\right)^{2n_{p}-1}$ (9)
@@ 2.2 Motzkin extension for forward scattering
The one-dimensional first-return probability with forward scattering is equation (2), encoding two processes: backward steps (probability $r$) and non-backward steps (probability 1r1-r), with the Motzkin triangle coefficients counting the combinatorial arrangements.
@@ 3.1 Physical origin
In bulk scattering without boundaries, Pfeifer and Chapman [13] proved that the Henyey–Greenstein phase function is closed under successive scattering: after $n$ scattering events, the angular distribution of the scattering cosine $\cos\theta=\hat{\Omega}_{0}\cdot\hat{\Omega}_{n}$ is itself a Henyey–Greenstein distribution with parameter $g^{n}$:
@@
$g_{\mathrm{eff}}^{(\mathrm{constrained})}=g^{n}\cdot\mathrm{BTF}(n,g)$ (13)
@@
$\mathrm{BTF}(n,g)=\frac{\int\cdots\int\prod_{i=1}^{n}P_{\mathrm{HG}}(\mu_{i};g)\times[\text{return constraint}]\,d\mu_{1}\cdots d\mu_{n}}{\int\cdots\int\prod_{i=1}^{n}P_{\mathrm{HG}}(\mu_{i};g)\,d\mu_{1}\cdots d\mu_{n}}$ (14)
@@ 3.2 Empirical discovery
For each $(g,n)$ pair, we extracted the BTF as the value needed to make the 1D formula reproduce the Monte Carlo 3D result:
$P_{3\mathrm{D}}^{(\mathrm{MC})}(n,g)=P_{\mathrm{1D}}^{\mathrm{marg}}(n,\,r_{b}(g,n))$ (15)
@@ 3.3 The Cauchy kernel
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\dfrac{n-2}{m_{x}(g)}\right)^{2}}$ (16)
with parameters:
$A(g)$ $=1-\frac{g(1+g)}{2}=\frac{(1-g)(2+g)}{2}\qquad\text{(amplitude)}$ (17)
$m_{x}(g)$ $=\frac{4g}{1-g}\qquad\text{(width)}$ (18)
The peak location $n_{0}=2$ corresponds to the minimum scattering order for first return.
Conjecture (Cauchy BTF).For Henyey–Greenstein scattering with anisotropy factor $g<2/3$, the Boundary Truncation Factor $\mathrm{BTF}(n,g)$ is numerically indistinguishable, within Monte Carlo precision, from a Cauchy kernel of the form (16) with parameters (17)–(18). At present this is supported by Monte Carlo evidence (Section 4) but lacks a rigorous derivation.
This parameterized form reproduces Monte Carlo-derived BTF values with mean deviation <<2% and cross-validated $R^{2}>0.999$ for $g\leq 2/3$.
@@
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation (16)). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible.
@@ 4.1 Simulation procedure
We follow the standard Monte Carlo procedure for photon transport [18]. First return occurs when $z$ first becomes negative [19]. Photons are initialized at the origin ($z=0^{+}$) with direction cosine $\mu_{0}=1$. At each scattering event, path length is sampled from $p(\ell)=e^{-\ell}$ (unit mean free path) and scattering angle from the Henyey–Greenstein phase function via inverse CDF. The scattering order $n$ is recorded at first occurrence of $z<0$. We simulated $10^{8}$ trajectories per $g$ value with 10 independent runs.
@@
$P_{\mathrm{HG}}(\mu;g)=\frac{1}{2}\,\frac{1-g^{2}}{(1+g^{2}-2g\mu)^{3/2}}\,,\qquad-1\leq\mu\leq 1$ (19)
where $g=\langle\mu\rangle\in[0,1)$ is the anisotropy parameter. The normalization $\int_{-1}^{1}P_{\mathrm{HG}}(\mu;g)\,d\mu=1$ is ensured by the prefactor $1/2$.
@@ 4.2 Fit quality
Error bounds. For $g<2/3$, root-mean-square deviation is below 2% across all $n$. For $g>2/3$, errors grow approximately as $\delta(g)\approx 0.05\cdot(g-2/3)/(1-g)$, reaching $\sim$15% at $g=0.9$. For $g>0.8$, the modified kernel (Section 7) is recommended.
5 Computational algorithm: normal incidence
We present the complete algorithm for mapping 3D Henyey–Greenstein scattering to 1D Motzkin combinatorics at normal incidence ($\mu_{\mathrm{inc}}=1$). The algorithm has four steps: (1) compute the single-scattering return probability; (2) determine the angular threshold; (3) evaluate the effective backscattering probability; (4) apply the Motzkin formula.
Step 1 (Single-scattering return probability): The two-step return probability anchors the entire framework:
$p_{r2}(g)=\int_{-1}^{0}\frac{\mu}{\mu-1}\,P_{\mathrm{HG}}(\mu,g)\,d\mu$ (20)
The geometric factor $\mu/(\mu-1)$ weights the phase function by the probability of returning to $z=0$ given exit direction $\mu$.
Step 2 (Angular threshold): The threshold $\mu_{b}(g)$ is the direction cosine separating “effective backward” from “effective forward” scattering in the 1D projection. Physically, $\mu_{b}$ partitions the angular phase space so that the 1D Motzkin counting matches the 3D return statistics. Solve the self-consistency condition:
$p_{r2}(g)=\tfrac{1}{2}\,F(-\mu_{b};\,g^{2})$ (21)
where $F(\mu;g)$ is the Henyey–Greenstein CDF (cumulative distribution function):
$F(\mu;g)=\int_{-1}^{\mu}P_{\mathrm{HG}}(\mu^{\prime};g)\,d\mu^{\prime}=\frac{1-g^{2}}{2g}\left[\frac{1}{\sqrt{1+g^{2}-2g\mu}}-\frac{1}{1+g}\right]$ (22)
with boundary values F(1;g)=0F(-1;g)=0 and $F(+1;g)=1$.
The parameter $g^{2}$ appears because after two scattering events the angular distribution has parameter $g^{2}$ (equation (12)).
Step 3 (Effective backscattering): In bulk scattering, the effective anisotropy after $n$ events is $g^{n}$ (standard convolution). Near a boundary, transverse filtering reduces this—captured by the BTF. The effective backscattering probability is
$r_{b}(g,n)=F\!\left(-\mu_{b}(g);\,g^{\,2\,+\,\mathrm{BTF}(n,g)\,(n-2)}\right)$ (23)
The exponent interpolates between $2$ at $n=2$ (where $\mathrm{BTF}=A(g)$ is maximal and the constraint is weakest) and $2$ as $n\to\infty$ (where $\mathrm{BTF}\to 0$ and boundary truncation dominates). The baseline of $2$ reflects the minimum scattering order for first return. This ansatz is empirically motivated: it produces the correct limits and matches Monte Carlo across the parameter space.